# Hypothetical Documents or Knowledge Leakage? Rethinking LLM-based Query Expansion

### v1

@@ Abstract
Query expansion methods powered by large language models (LLMs) have demonstrated effectiveness in zero-shot retrieval tasks. These methods assume that LLMs can generate hypothetical documents that, when incorporated into a query vector, enhance the retrieval of real evidence. However, we challenge this assumption by investigating whether knowledge leakage in benchmarks contributes to the observed performance gains. Using fact verification as a testbed, we analyzed whether the generated documents contained information entailed by ground truth evidence and assessed their impact on performance. Our findings indicate that performance improvements occurred consistently only for claims whose generated documents included sentences entailed by ground truth evidence. This suggests that knowledge leakage may be present in these benchmarks, inflating the perceived performance of LLM-based query expansion methods, particularly in real-world scenarios that require retrieving niche or novel knowledge.
@@ 1 Introduction
In this paper, we challenge the underlying assumption, as illustrated in Figure 1: Do LLMs truly generate hypothetical documents, or are they merely reproducing what they already know? LLMs are extensively pretrained on vast corpora, primarily collected from the web. As common retrieval targets, such as Wikipedia and web documents, are often included in these pretraining corpora Groeneveld et al. (2024); Touvron et al. (2023); Du et al. (2022); Brown et al. (2020), many available LLMs may already contain knowledge relevant to a given query and retrieval targets, a phenomenon we refer to as knowledge leakage. If knowledge leakage occurs, it could lead to an overestimation of the effectiveness of QE methods in real-world scenarios, where generating documents for unknown knowledge is crucial for recent or niche queries.
To understand whether knowledge leakage exists and how it influences benchmark performance, this study examines LLM-generated documents using fact verification as a testbed. We analyze whether these documents contain sentences entailed by gold evidence and assess their impact on performance. Across experiments involving three benchmarks and seven LLMs, we observed a consistent trend: query expansion methods were effective only when LLM-generated documents included sentences entailed by gold evidence. This result implies that knowledge leakage within these benchmarks may artificially boost the observed effectiveness of LLM-based query expansion approaches.
@@ 2 Related Works
Previous research has investigated various forms of data leakage in LLMs Kandpal et al. (2023); Samuel et al. (2025); Deng et al. (2024); Xu et al. (2024a). One study used perplexity to detect potential data leakage, uncovering substantial instances of training or even test set misuse Xu et al. (2024b). Deng et al. (2023) further examined data contamination by predicting masked tokens in test sets and found that GPT 3.5 could reconstruct missing portions of MMLU Hendrycks et al. (2020) test instances with 57% accuracy. Other research has explored the boundaries of LLM knowledge Yin et al. (2023); Dong et al. (2024); Burns et al. (2022); Kadavath et al. (2022), including a refusal-aware instruction tuning method that trains LLMs to reject uncertain questions Zhang et al. (2024)—where an LLM is deemed uncertain if its generated response does not match the ground truth. Another study leveraged response consistency to estimate an LLM’s confidence in its knowledge Cheng et al. (2024). In this work, we apply NLI MacCartney (2009) to LLM-generated documents, referencing gold evidence, to examine what an LLM knows.
@@ 3.1 Task and Dataset
The fact-verification task consists of two sub-tasks: evidence retrieval and verdict prediction. In evidence retrieval, a retrieval method $R(\cdot)$ identifies an evidence set $\tilde{E}=\{\tilde{e_{1}},\cdots,\tilde{e_{k}}\}$ from a knowledge store $K$ (e.g., Wikipedia), used to verify $c$. The performance of $R$ is evaluated by comparing $\tilde{E}$ with $E=\{e_{1},\cdots,e_{l}\}$, the gold evidence set. Verdict prediction then determine the veracity label of $c$ based on $\tilde{E}$.
@@
Table 2: Fact verification performance using baseline retrievers and LLM-based query expansion methods, with the number of retrieved evidence set to five ($k=5$). We report the average performance of query expansion methods along with standard errors, obtained by repeating the generations eight times.
We chose fact verification as the target task to test our hypothesis for two reasons. First, in real-world fact-checking scenarios Miranda et al. (2019); Nakov et al. (2021), retrieving evidence about niche or novel knowledge is crucial. If QE is effective only when the relevant knowledge has been seen during language model pretraining, its practical usefulness could be limited. Second, verdict prediction is a classification task, thus facilitating clearer evaluation of how QE influences final outcomes compared to generation-based tasks, such as factual QA Joshi et al. (2017); Kwiatkowski et al. (2019).
We employ three datasets that provide annotated evidence and corresponding veracity labels, along with an external knowledge store: FEVER Thorne et al. (2018), SciFact Wadden et al. (2020), and AVeriTeC Schlichtkrull et al. (2023). While FEVER and SciFact contain verbatim extractions from $K$ as gold evidence (i.e., $E$), AVeriTeC uses human-written evidence referencing $K$ to verify claims. The three datasets cover diverse knowledge domains, as summarized in Table 1.
3.2 LLM-based Query Expansion
We evaluate two representative LLM-based QE methods, which expand queries by prompting an LLM.
Query2doc Wang et al. (2023)
$q^{+}=concat(q\times n,d)$ (1)
HyDE Gao et al. (2023)
employs an LLM to generate hypothetical documents $[d_{1},...,d_{N}]$ to answer a query $q$. A dense retriever $g(\cdot)$ encodes $q$ and each $d_{k}$ separately, and their encoded embeddings are averaged to form the query vector $v_{q^{+}}$ (Equation 2).
$v_{q^{+}}=\frac{1}{N+1}\sum_{k=1}^{N}[g(d_{k})+g(q)]$ (2)
@@ 3.2 LLM-based Query Expansion
Table 3: Proportion of expanded queries containing sentences entailed by ground truth evidence across different LLM backbones.
3.3 Matching Method
[…]
We use the sentence segmentation module provided by spaCy [^2] https://huggingface.co/spacy/en_core_web_lg, and employ GPT-4o-mini for NLI (Figure A3). To filter out claim reproductions, we apply ROUGE-2 Lin (2004) with a threshold of 0.95, based on manual inspection.
4 Experimental Results
We conducted evaluation experiments on three fact verification benchmarks. Each LLM-based generation was repeated eight times, and we report the average performance with the standard error.
@@ 4 Experimental Results
To assess the effectiveness of Query2doc and HyDE, we compared their performance against BM25 and Contriever that use $c$ as query, respectively, as baseline retrievers. For evidence retrieval, we used Recall@$k$ and NDCG@$k$ ($k=5$) as evaluation metrics Manning (2009) on the FEVER and SciFact datasets, where both the ground-truth evidence $E$ and retrieved evidence $\tilde{E}$ come from the knowledge store $K$. In contrast, $\tilde{E}$ in AveriTeC consists of human-written evidence rather than extracts from $K$. Therefore, following previous studies Schlichtkrull et al. (2023); Chen et al. (2022), we applied the Hungarian algorithm Kuhn (1955) with METEOR Banerjee and Lavie (2005) and BERTScore Zhang et al. (2020) on the top five retrieved sentences, computing token-level and embedding-level similarity, respectively, based on a binary assignment between generated and reference sequences. For verdict prediction, we used GPT-4o-mini with the five retrieved evidence and evaluated performance using macro F1. Evaluation details are provided in Appendix B.
@@
Table 3 presents the proportion of matched claims across the three datasets and seven LLMs. In most cases, more than 40% of the claims were matched, with a few exceptions. The lowest proportion (27.6%) was observed for SciFact when claims were expanded using HyDE with Gemini-1.5-flash—still a notable fraction. The highest proportion (83.5%) was observed for FEVER when using HyDE with GPT-4o-mini. Several examples of LLM-generated documents and gold evidence are provided in Table A8.
@@
Table 4 presents fact-verification performance based on whether LLM-generated documents contained sentences entailed by gold evidence, focusing on GPT-4o-mini. We observed a consistent trend across the three datasets and two expansion methods: matched claims (where LLM-generated documents contain sentences entailed by gold evidence) achieved significantly better performance than both all and unmatched claims, with statistical significance at p$<$0.001. Moreover, with a few exceptions, performance on unmatched claims was lower than that of the respective baseline methods without query expansion—BM25 for Query2doc and Contriever for HyDE. This trend was consistent across seven different LLMs and for $k=10$, as shown in Table A2 and Table A3.
@@ 5 Discussion
Our findings provide empirical support for the effectiveness of LLM-based query expansion in zero-shot retrieval tasks Gao et al. (2023); Jagerman et al. (2023); Lei et al. (2024); Wang et al. (2024a), particularly on fact verification benchmarks. They also highlights the crucial role of retrieving suitable evidence for claim verification Chen et al. (2024a); Guo et al. (2022), suggesting a need for further research into evidence retrieval techniques, such as claim decomposition Chen et al. (2022); Fan et al. (2020).
@@
This finding suggests that the success of hypothetical document generation, observed in Table 2 and prior studies Gao et al. (2023); Wang et al. (2023); Yoon et al. (2024), may be largely attributable to LLM’s internal knowledge encompassing benchmark knowledge sources. Furthermore, these results suggest that LLM-based query expansion may be limited in real-world scenarios requiring the retrieval of niche or novel knowledge, such as fact verification in the wild. Future advances in query expansion could incorporate external knowledge sources to address these limitations, as demonstrated in recent work Lei et al. (2024).
@@ 6 Conclusion
This study examined the impact of two widely used LLM-based query expansion methods on fact verification. By applying NLI to the LLM-generated documents, we identified a consistent trend suggesting that knowledge leakage may be present in these benchmarks, inflating the perceived performance of these methods, particularly in real-world scenarios involving niche or novel knowledge. These findings highlight the need for future research on LLM-based query expansion methods that can effectively handle unknown queries, as well as the development of evaluation frameworks that more accurately reflect real-world settings.
@@
This study analyzed LLM behaviors to identify potential indicators of memorization and their impact on performance but did not investigate whether training on specific knowledge directly led to the generation of corresponding knowledge. Therefore, we cannot establish a causal relationship between data leakage and generation in response to query expansion prompts. Future research could further validate our methodology and findings by conducting experiments with synthetic data Liu et al. (2024) or evaluating LLMs on genuinely novel knowledge Kasai et al. (2024). To support the validity of the NLI-based automatic evaluation, we conducted a manual validation on sampled data and observed a consistent trend (Table A4) with that of the automatic method (Table 4).
@@
We applied an NLI-based algorithm to investigate whether documents generated by LLM-based query expansion methods share underlying knowledge with gold evidence in fact verification benchmarks. Our findings suggest that knowledge leakage may exist; namely, LLMs may have been exposed to information related to the benchmarks’ gold evidence during training and subsequently reproduced it in response to query-expansion prompts. This observation has important implications for both fact verification and broader benchmark-oriented NLP research, as it suggests that benchmark performance may be artificially inflated. Consequently, more trustworthy evaluation frameworks are needed to accurately reflect real-world scenarios.
Despite these observations, some caution is warranted when interpreting these findings. First, this study did not provide definitive evidence of knowledge leakage. As outlined in the Limitations, our research design cannot examine the causal relationship between model training data and the generated outputs. Future research could address this gap by conducting (continued) pretraining experiments using synthetic data or genuinely novel knowledge. Second, because our analysis focuses on three datasets within the fact-verification domain, the findings are limited in scope. Further experiments are necessary to determine whether these findings generalize to other tasks, such as factual QA Joshi et al. (2017); Kwiatkowski et al. (2019). While the methodology presented here could be adapted to those settings, there remains a risk that the NLI model itself may introduce inaccuracies or biases toward particular labels. We conducted manual validation of the NLI results to mitigate these risks. Finally, we used ChatGPT to proofread portions of this manuscript.
@@ Acknowledgments
This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the Graduate School of Metaverse Convergence support (IITP-2025-RS-2024-00430997) and Innovative Human Resource Development for Local Intellectualization (IITP-2025-RS-2022-00156360) programs, supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation). Kunwoo Park is the corresponding author.
@@ Appendix D Computing Environment
We ran experiments using two machines. The first machine has four Nvidia RTX A6000 GPUs (48GB per GPU) and 256GB RAM. The second machine has two Nvidia H100 GPUs (80GB per GPU) and 480 RAM. The experiments were conducted in a computing environment with the following configuration: Python 3.11.10, PyTorch 2.3.1, Transformers 4.43.4, vLLM 0.5.3, pyrec-eval 0.5, Faiss 1.8, Pyserini 0.40.0, NLTK 3.9.1, bert-score 0.3.13, rouge-score 0.1.2.

### v2

@@ Abstract
Query expansion methods powered by large language models (LLMs) have demonstrated effectiveness in zero-shot retrieval tasks. These methods assume that LLMs can generate hypothetical documents that, when incorporated into a query vector, enhance the retrieval of real evidence. However, we challenge this assumption by investigating whether knowledge leakage in benchmarks contributes to the observed performance gains. Using fact verification as a testbed, we analyze whether the generated documents contain information entailed by ground-truth evidence and assess their impact on performance. Our findings indicate that, on average, performance improvements consistently occurred for claims whose generated documents included sentences entailed by gold evidence. This suggests that knowledge leakage may be present in fact-verification benchmarks, potentially inflating the perceived performance of LLM-based query expansion methods.
@@ 1 Introduction
In this paper, we challenge the underlying assumption, as illustrated by a counterexample in Figure 1: Do LLMs truly generate hypothetical documents, or are they merely reproducing what they already know? LLMs are extensively pretrained on vast corpora, primarily collected from the web. As common retrieval targets, such as Wikipedia and web documents, are often included in these pretraining corpora Groeneveld et al. (2024); Touvron et al. (2023); Du et al. (2022); Brown et al. (2020), many available LLMs may already contain knowledge relevant to a given query and retrieval targets in benchmark datasets, a phenomenon we refer to as knowledge leakage. If knowledge leakage occurs and an LLM reproduces what it already knows about a given query, its generation cannot be seen as hypothetical documents. Furthermore, assessing QE performance on a benchmark where leakage occurs could lead to overestimation, as retrieving verbatim or semantically equivalent sentences is a much more trivial task than predicting the relevance between a query and document when the relevance signal is not explicit.
To investigate whether knowledge leakage exists in popular benchmarks and how it affects downstream performance, this study uses fact verification as a testbed. We analyze whether LLM-generated documents contain sentences entailed by gold evidence and assess their impact on performance. Across experiments spanning three benchmarks and seven LLMs, we observed a consistent trend: QE methods were effective, on average, only when the generated documents included sentences entailed by gold evidence. These findings suggest that knowledge leakage may be present in these benchmarks, potentially inflating the perceived effectiveness of LLM-based QE methods.
@@ 2 Related Works
Previous research has investigated various forms of data leakage in LLMs Kandpal et al. (2023); Samuel et al. (2025); Deng et al. (2024); Xu et al. (2024a). One study used perplexity to detect potential data leakage, uncovering substantial instances of training or even test set misuse Xu et al. (2024b). Deng et al. (2023) further examined data contamination by predicting masked tokens in test sets and found that GPT 3.5 could reconstruct missing portions of MMLU Hendrycks et al. (2020) test instances with 57% accuracy. Other research has explored the boundaries of LLM knowledge Yin et al. (2023); Dong et al. (2024); Burns et al. (2022); Kadavath et al. (2022); Mallen et al. (2023), including a refusal-aware instruction tuning method that trains LLMs to reject uncertain questions Zhang et al. (2024)—where an LLM is deemed uncertain if its generated response does not match the ground truth. Another study leveraged response consistency to estimate an LLM’s confidence in its knowledge Cheng et al. (2024). In this work, we apply NLI MacCartney (2009) to LLM-generated documents, referencing gold evidence, to examine what an LLM knows.
@@ 3.1 Task and Dataset
The fact-verification task consists of two sub-tasks: evidence retrieval and verdict prediction. In evidence retrieval, a retrieval method $R(\cdot)$ identifies an evidence set $\tilde{E}=\{\tilde{e_{1}},\cdots,\tilde{e_{k}}\}$ from a knowledge store $K$ (e.g., Wikipedia), used to verify $c$. The performance of $R$ is evaluated by comparing $\tilde{E}$ with $E=\{e_{1},\cdots,e_{l}\}$, the gold evidence set. Verdict prediction then determines the veracity label of $c$ based on $\tilde{E}$.
We choose fact verification as the target task to test our hypothesis for two main reasons. First, in real-world fact-checking scenarios Miranda et al. (2019); Nakov et al. (2021), retrieving evidence about niche or novel knowledge is crucial. If QE is effective only when the relevant knowledge has been seen during language model pretraining, its practical usefulness may be limited. Second, verdict prediction is a classification task, which enables clearer evaluation of how QE influences end performance compared to generation-based tasks, such as factual QA Joshi et al. (2017); Kwiatkowski et al. (2019).
We employ three datasets that provide annotated evidence and corresponding veracity labels, along with an external knowledge store: FEVER Thorne et al. (2018), SciFact Wadden et al. (2020), and AVeriTeC Schlichtkrull et al. (2023). While FEVER and SciFact contain verbatim extractions from $K$ as gold evidence (i.e., $E$), AVeriTeC uses human-written evidence referencing $K$ to verify claims. These three datasets cover diverse knowledge domains, as summarized in Table 1.
3.2 LLM-based Query Expansion
We consider two LLM-based QE methods.
Query2doc Wang et al. (2023)
$q^{+}=concat(q\times n,d)$ (1)
HyDE Gao et al. (2023)
employs an LLM to generate hypothetical documents $[d_{1},...,d_{N}]$ in response to a query $q$. A dense retriever $g(\cdot)$ encodes $q$ and each $d_{k}$ separately, and their encoded embeddings are averaged to form the query vector $v_{q^{+}}$ (Equation 2).
$v_{q^{+}}=\frac{1}{N+1}\sum_{k=1}^{N}[g(d_{k})+g(q)]$ (2)
3.3 Matching Algorithm
We present an algorithm to determine whether an LLM-generated document $d$ for a claim $c$ contains a sentence that is verbatim or semantically equivalent to the gold evidence $E=\{e_{1},\cdots,e_{m}\}$. If such a sentence exists, it may indicate that the backbone LLM was exposed to knowledge related to the corresponding evidence during training. To automatically identify such claims, we rely on natural language inference (NLI) assigning each claim $c$ to one of two conditions: matched ($M$) or unmatched ($\neg M$). The algorithm consists of three steps.
  • (1)
    Sentence Segmentation: Segment $d$ into sentences and remove reproductions of $c$ to construct $S=\{s_{1},\cdots,s_{n}\}$.
  • (2)
[…]
  • (3)
    Label Aggregation: Aggregate all labels $\{l_{(1,1)},\cdots,l_{(i,j)},\cdots,l_{(m,n)}\}$ into a single label $l$. If there exists $l_{(i,j)}$ labeled as entailment, assign matched, otherwise assign unmatched.
We use the sentence segmentation module provided by spaCy [^2] https://huggingface.co/spacy/en_core_web_lg, and employ GPT-4o-mini for NLI (Figure A3). To filter out sentences of reproduced claims in LLM responses, we apply ROUGE-2 Lin (2004) with a threshold of 0.95, based on manual inspection.
@@ 4 Experimental Results
Table 2: Fact verification performance using baseline retrievers and LLM-based query expansion methods, with the number of retrieved evidence set to five ($k=5$).
@@
Table 3: Proportion of expanded queries containing sentences entailed by ground truth evidence.
We conducted evaluation experiments on three fact-verification benchmarks. Based on eight repetitions of LLM generation, we report the average performance along with the standard error.
@@
To assess the effectiveness of LLM-based QE methods, we compared their performance against BM25 and Contriever that use $c$ as query, respectively, as baseline retrievers. For evidence retrieval, we used Recall@$k$ and NDCG@$k$ ($k=5$) as evaluation metrics Manning (2009) on the FEVER and SciFact datasets, where both the ground-truth evidence $E$ and retrieved evidence $\tilde{E}$ come from the knowledge store $K$. In contrast, $\tilde{E}$ in AveriTeC consists of human-written evidence rather than extracts from $K$. Therefore, following previous studies Schlichtkrull et al. (2023); Chen et al. (2022), we applied the Hungarian algorithm Kuhn (1955) with METEOR Banerjee and Lavie (2005) and BERTScore Zhang et al. (2020) on the top five retrieved sentences, computing token-level and embedding-level similarity, respectively, based on a binary assignment between generated and reference sequences. For verdict prediction, we used GPT-4o-mini with the five retrieved evidence and evaluated performance using macro F1. Evaluation details are provided in Appendix B.
@@
Table 3 presents the proportion of matched claims observed for the three datasets by using seven different LLMs. In most cases, more than 40% of the claims were matched, with a few exceptions. The lowest proportion (27.6%) was observed for SciFact when claims were expanded using HyDE with Gemini-1.5-flash—still a notable fraction. The highest proportion (83.5%) was observed for FEVER when using HyDE with GPT-4o-mini. Several examples of LLM-generated documents and gold evidence are provided in Table A8.
@@
Table 4 presents fact-verification performance based on whether LLM-generated documents contained sentences entailed by gold evidence, focusing on GPT-4o-mini as the QE model. We observed a consistent trend across the three datasets and two expansion methods: performance measured on matched claims (where LLM-generated documents contain sentences entailed by gold evidence) was significantly higher than that on both all and unmatched claims, with statistical significance at p$<$0.001. Moreover, with a few exceptions, performance on unmatched claims was lower than that of the corresponding baseline methods without query expansion—BM25 for Query2doc and Contriever for HyDE. These trends were consistent across seven different LLMs and for $k=10$, as shown in Table A2 and Table A3.
@@ 5 Discussion
Our findings support the effectiveness of LLM-based query expansion in zero-shot retrieval tasks Gao et al. (2023); Jagerman et al. (2023); Lei et al. (2024); Wang et al. (2024a), particularly on fact-verification benchmarks. The increased verification performance through query expansion further highlights the crucial role of retrieving suitable evidence for claim verification Chen et al. (2024a), suggesting a need for further research into evidence retrieval methods, such as claim decomposition Chen et al. (2022); Fan et al. (2020).
@@
This finding suggests that the success of hypothetical document generation may be largely attributable to the LLM’s internal knowledge encompassing benchmark knowledge sources. Furthermore, these results indicate that LLM-based query expansion may be limited in real-world scenarios that require retrieving niche or novel knowledge. Future research could address these limitation by incorporating external knowledge sources into the query expansion process, as demonstrated in a recent study Lei et al. (2024).
@@ 6 Conclusion
This study examined the leakage hypothesis for query expansion, questioning whether an LLM generates truly hypothetical documents or merely reproduces what was exposed during the LLM’s training phase. Through experiments on three widely used fact-verification benchmarks, we observed a consistent trend: on average, retrieval and verification performance increased only when the expanded query was entailed by gold evidence. This finding suggests that knowledge leakage may be present in these benchmarks, artificially inflating the perceived effectiveness of LLM-based query expansion. Since verifying claims involving niche or novel knowledge is often essential in real-world scenarios, such leakage can hinder accurate evaluation and, consequently, impede method development. These results highlight the need for future research on LLM-based query expansion methods capable of handling unknown queries, as well as the creation of evaluation benchmarks that more accurately reflect real-world conditions involving previously unseen knowledge.
@@
While our analysis provides the evidence that certain LLM behaviors can signal memorization and that these behaviors correlate with downstream performance gains, we acknowledge three areas where further validation is needed. (1) Causal mechanism: Our experiments focus on LLM behaviors rather than tracing whether specific training data directly results in the generation of associated knowledge. Consequently, we do not claim a causal link between data leakage and generation in response to expansion prompts. Establishing such a connection remains an open challenge and would benefit from controlled studies that expose models to synthetic or novel knowledge Liu et al. (2024); Kasai et al. (2024). (2) NLI-based matching: To corroborate the NLI-based automatic assessment, we manually annotated a representative sample and observed a consistent trend (Table A4) with the automatic results (Table 4). Adopting a stronger NLI model would further strengthen our conclusions. (3) Generalizability: Our analysis focused on three datasets within the fact-verification domain, and hence the findings are limited in scope. Further experiments are necessary to determine whether these findings generalize to other tasks, such as factual QA Joshi et al. (2017); Kwiatkowski et al. (2019). The methodology presented here could be adapted to those settings.
@@
Our findings suggest that knowledge leakage may exist in three popular fact-verification benchmarks; namely, LLMs may have been exposed to information related to the benchmarks’ gold evidence during pretraining and subsequently reproduced it in response to query-expansion prompts. This observation has important implications for both fact verification and broader benchmark-oriented NLP research, as the performance of LLM-based methods may be artificially inflated due to the opaqueness of the pretraining stage. Consequently, more trustworthy evaluation frameworks and transparent LLM development practices, such as those exemplified by Groeneveld et al. (2024) and Gao et al. (2020), are encouraged to better reflect real-world performance. We used ChatGPT to proofread portions of this manuscript.
@@ Acknowledgments
This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the Innovative Human Resource Development for Local Intellectualization (IITP-2025-RS-2022-00156360) and Graduate School of Metaverse Convergence support (IITP-2025-RS-2024-00430997) programs, supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation). Kunwoo Park is the corresponding author.
@@ Appendix D Computing Environment
We ran experiments using two machines. The first machine has four Nvidia RTX A6000 GPUs (48GB per GPU) and 256GB RAM. The second machine has two Nvidia H100 GPUs (80GB per GPU) and 480GB RAM. The experiments were conducted in a computing environment with the following configuration: Python 3.11.10, PyTorch 2.3.1, Transformers 4.43.4, vLLM 0.5.3, pyrec-eval 0.5, Faiss 1.8, Pyserini 0.40.0, NLTK 3.9.1, bert-score 0.3.13, rouge-score 0.1.2.